OpenStudy (anonymous):

The fastest remote-controlled car can reach 55.09 mph. If its battery lasts 6 hours and 10 minutes, which of these is the approximate maximum distance the car can go on one battery charge? 332.38 miles 348.9 miles 340.54 miles 339.72 miles

OpenStudy (studygurl14):

Hey @ashsek you wanna take this one?

OpenStudy (anonymous):

try it yourself any guss

OpenStudy (anonymous):

ok

OpenStudy (studygurl14):

@judtpleasegahd yea I've shown you how to do several similar ones, so I bet you can do this. :)

OpenStudy (anonymous):

You stupid

OpenStudy (studygurl14):

@fuzzyowls please do not spam questions with comments like that

OpenStudy (anonymous):

is the answer 339.72

OpenStudy (anonymous):

Can anybody help me with 2 questions?

OpenStudy (studygurl14):

lol yea @ashsek but i wanted @judtpleasegahd to do it....

OpenStudy (anonymous):

ya let her do this alone

OpenStudy (studygurl14):

@fuzzyowls do not hijack other people's questions with your own. instead, tag people to your question by using the @ symbol

OpenStudy (anonymous):

@fuzzyowls

OpenStudy (anonymous):

1.Name five fractions equivalent to 4/12 2. Name five fractions that are near, but not equivalent to 4/12

OpenStudy (studygurl14):

@fuzzyowls if you do not stop, I will have to report you.

OpenStudy (anonymous):

lol ok back to what i was doingg

OpenStudy (studygurl14):

yep. So, how do get the answer of 339.72 with the information given? @judtpleasegahd

OpenStudy (anonymous):

well its not c,... D

OpenStudy (anonymous):

ok divde first 6 and 10

OpenStudy (studygurl14):

um...no. If they say the car goes 55.09mph, and lasts 6hours and 10 minutes, what do you do to find how many miles it can go?

OpenStudy (anonymous):

divide

OpenStudy (anonymous):

try to use the formulae \[speed = \frac{ Distance }{time }\]

OpenStudy (studygurl14):

using the info @ashsek gave: "55.09 mph" "6 hours and 10 minutes: plug in and solve for distance \(\large 55.09=\frac{distance}{6\frac{10}{60}}\)